OpenStudy (anonymous):

(x)/(4) + (2x-1)/(3) - (x+2)/(6) = ?

OpenStudy (pfenn1):

\[\left( x \over 4 \right)+\left( 2x-1 \over 3 \right)-\left( x+2 \over 6 \right)=\] what is the common denominator?

OpenStudy (anonymous):

Um, 12?

OpenStudy (pfenn1):

Right. Do you know how to get all the terms to have that same denominator?

OpenStudy (anonymous):

Nuhuh.

OpenStudy (pfenn1):

Ok. For the first term, we can multiply the term by 3/3\[\left( x \over 4 \right)\left( 3\over 3 \right)=\left( 3x \over 12 \right)\]

OpenStudy (anonymous):

So you multiply it by whatever number you can times it by to get 12?

OpenStudy (pfenn1):

The second term we can multiply by 4/4\[\left( 2x-1 \over 3 \right)\left( 4 \over 4 \right)=\left( 8x-4 \over 12 \right)\]

OpenStudy (pfenn1):

Correct. So what would the last term be?

OpenStudy (anonymous):

(2x+4)/(12)

OpenStudy (pfenn1):

yep. Now put all the terms back into the initial equation\[\left( 3x \over 12 \right)+\left( 8x-4 \over 12 \right)-\left( 2x+4 \over 12 \right)=\]

OpenStudy (pfenn1):

Simplify

OpenStudy (anonymous):

(9x)/12

OpenStudy (pfenn1):

Hmmmm. I got \[9x-8 \over 12\]

OpenStudy (anonymous):

how. 4+(-4)=0

OpenStudy (pfenn1):

There is a negative sign in front of that last term so 4-(-4)=-8

OpenStudy (pfenn1):

or rather -4-(4)=-8

OpenStudy (anonymous):

ohh, okay. I see now. So can you help wit this one, cause i dont understand how to do common denominators with variables. (1)/(a) - (6)/(bc) + (3)/(ac) =

OpenStudy (pfenn1):

Sure.

OpenStudy (anonymous):

Thank you.

OpenStudy (pfenn1):

\[\left( 1 \over a \right)-\left( 6 \over bc \right)+\left( 3 \over ac \right)=\] what would be the common denominator?

OpenStudy (anonymous):

im not sure, i dont understand. abc?

OpenStudy (pfenn1):

yes it is abc. so to get all the terms to have the same common denominator abc, we would multiply the first term by bc/bc, the second term by a/a, and the third term by b/b. \[\left( bc \over abc \right)-\left( 6a \over abc \right)+\left( 3b \over abc \right)=\]

OpenStudy (anonymous):

Oh ok, i see now.

OpenStudy (pfenn1):

Simplify and get\[\left( bc-6a-3b \over abc \right)=\left( (c-3)b-6a \over abc \right)\]

OpenStudy (pfenn1):

I don't know if you need that last step....you could leave it in either form

OpenStudy (anonymous):

Ok, thank you so much. I understand much better now. Can you help me with two more? Or i can make a new thread and get someone else.

OpenStudy (pfenn1):

Sure, but I may have to start dinner soon.

OpenStudy (anonymous):

I understand that, ill find someone else. thank you so much for helping me.

OpenStudy (pfenn1):

Good luck!